OpenStudy (anonymous):

Which of the following is NOT equal to the others? A) 12 - x B) x - 12 C) -(x - 12)

myininaya (myininaya):

B

myininaya (myininaya):

12-x=-x+12=-(x-12)

OpenStudy (anonymous):

Thnks